OpenStudy (anonymous):

The length of two sides of a right triangle are given. Find the length of the third side. Round to the nearest tenth. 28in and 15in A.37.5 B.23.6 C.29.6 D.31.8

OpenStudy (anonymous):



OpenStudy (anonymous):

Assuming you are given "two legs" of the right traingle and question is asking to find the hypotenuse.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

to find the length og the hypotenuse, you can use: a² + b² = c² and c is the hypotenuse. so we know a = 28, and b = 15 28² + 15² = c² 784 + 225 = c² 1009 = c² \[\sqrt{1009} = \sqrt{c²}\] 31.8 = c